OpenStudy (anonymous):

How can i do this??? a rectangle is 8 feet ling and 6 feet wide. if each side is increased by the same amount, the area of the new rectangle is 72 square feet more than the area of the original rectangle . find the length and width of the new rectangle in feet..

OpenStudy (anonymous):

(8+a)(6+a)=8*6 +72 now solve for a

OpenStudy (diyadiya):

Area of given rectangle = length x breadth =6 x 8=48 ft^2 Area of New rectangle =(8+x)(6+x)= 48 +72

OpenStudy (mathteacher1729):

area = length times width A = LW The original rectangle had area 8ft * 6ft = 48 ft^2 The new rectangle is 72ft^2 more so ... (8+x)*(6+x) = 48 ft^2 + 72 ft^2 You must solve for x. :)

OpenStudy (anonymous):

12ft n 10ft

OpenStudy (diyadiya):

(8+x)(6+x)=120 48+8x+6x+x^2=120 x^2+14x+48=120 x^2+14x-72=0

OpenStudy (anonymous):

Thanks guys!!! :)